# Frozen Transformers in Language Models Are Effective Visual Encoder Layers

### v1

@@ 6.3 Information Filtering Hypothesis on Other Tasks
The previous sections mainly discuss our information filtering hypothesis in terms of image classification. Meanwhile, we also discover supportive evidence of our hypothesis on various other tasks. Due to space limits, this section investigates action recognition as an example, and Sec. A.2 covers additional tasks including point cloud classification, 2D VQA, and 3D VQA.
@@
We mainly analyze the information filtering hypothesis in action recognition qualitatively because the ground truth of “relevant regions” is difficult to quantify for this task. In practice, we follow a similar procedure in Sec. 6.1 and visualize the activation of video tokens in Fig. 5. At a low activation threshold, we notice that the video tokens from ViT-S-LLaMA better capture the foreground areas of hands and manipulated objects than ViT-S. With the video tokens in VideoMAE activated both spatially and temporally, we further increase the threshold to demonstrate its ability to select informative frames. As in the “high threshold” row of Fig. 5, ViT-S-LLaMA more accurately focuses on the middle frames with actual human-object interaction. Therefore, we conclude that the informative video tokens are indeed distinguished and augmented in action recognition, which aligns with the information filtering hypothesis.
6.4 Discussion and Limitations of the Hypothesis
Perspective of usable information.   We supplement the opinions from Xu et al. (2020) that greatly inspire our investigation of the information filtering hypothesis. Xu et al. (2020) proposes that a well-trained neural network layer can be considered as a decipher adding usable information into the features and enabling subsequent modules to better infer the latent information. In our information filtering hypothesis, the incorporated modules are indeed acting as deciphers that enlarge the contribution of usable tokens and benefit downstream predictions.
7 Conclusions
Discussion and limitations.   We have validated the capability of pre-trained, frozen language transformers across a wide spectrum of visual tasks. It is important to note that our goal is to methodically explore this under-investigated problem. Therefore, our experiments are designed to maximize the diversity of tasks under fair comparisons with well-established or competitive baselines, rather than striving for state-of-the-art performance for all tasks, which is also constrained by our computational resources. We leave scaling up the experiments to state-of-the-art levels for all tasks as interesting future work. Meanwhile, we also notice that our information filtering hypothesis has not covered several intriguing questions, e.g., how to quantify the functions of different layers and analyze how the training process facilitates the visual tokens features to cooperate with the language transformer, which are also meaningful directions.
@@ 7 Conclusions
Conclusions.   In this work, we explore the unexpected capability of large language models (LLMs) as encoders for visual tasks, a significant departure from their conventional text-based applications. By seamlessly integrating a frozen transformer block from pre-trained LLMs into visual encoders, we observe consistent performance enhancements across diverse visual challenges from 2D image and video classification, 3D point cloud recognition, non-semantic motion forecasting, as well as 2D and 3D vision-language tasks. This phenomenon, underpinned by our proposed information filtering hypothesis, highlights the inherent adaptability and versatility of LLMs for more general representation learning. We hope that our insights will catalyze further exploration into the uncharted fields of LLM applications and foster innovative strategies to harness their potential in novel ways.
Acknowledgement.   We thank Baifeng Shi and Shoufa Chen for their valuable discussion and suggestions on implementation and model training.
A Appendix for the Information Filtering Hypothesis
@@ A.1 Detailed Derivation of the Hypothesis
which is identical to the main paper (Eqn. 3), we further separate the visual tokens into the subset of informative tokens $V_{i}$ and uninformative tokens $V_{u}$. This changes Eqn. A into,
@@
Using the same notation, our observation on the feature activation can be stated as: the attention weights for informative tokens $\{w_{i},i\in V_{i}\}$ are still noisy after incorporating the frozen LLM transformer, while the final visual tokens shown as below have emergent concentration on target regions.
$z_{L}^{2}[i]=\mathbf{F}_{L}^{2}\cdot\mathbf{F}_{LM}^{F}\cdot\mathbf{F}_{LM}^{A}(z_{L}^{1}[i]),\ \text{where}\ i\in V_{i}$ (C)
Combining Eqn. B and Eqn. C inspires us to express our hypothesis in terms of the connection between visual and $\mathtt{CLS}$ token with Eqn. D below, where the added modules $\mathbf{F}_{L}^{2}\cdot\mathbf{F}_{LM}^{F}\cdot\mathbf{F}_{LM}^{A}$ augment the informative tokens $V_{i}$ and lead to better prediction:
@@ A.2 Information Filtering Hypothesis on Other Tasks
This section provides the evidence for our information filtering hypothesis in other tasks, supplementing the discussion on image classification (Sec. 6.1) and action recognition (Sec. 6.3). Specifically, we observe that the frozen LLM transformer selects and amplifies information tokens in point cloud recognition, 2D visual question answering (VQA), and 3D VQA. The tasks of motion forecasting and image-text retrieval are not illustrated because it is more abstract to intuitively define their “informative” tokens. For the investigated tasks, we mainly analyze qualitatively because the ground truth for relevant regions is ambiguous on such tasks, unlike the segmentation masks for image classification (Sec. 6.2).
@@
Point cloud recognition.   We visualize the activation of the point tokens in point cloud classification before and after adding the frozen LLM transformer in Fig. A. In the examples of chairs and desks, we observe that “PointBERT+LLaMA” concentrates less on the background (chairs, indicated with red arrows) and more on the actual object surfaces (desks, indicated with red arrows). This demonstrates that the frozen LLM transformer learns to focus on the informative points, which is consistent with our hypothesis.
@@
Figure A: Visualization of feature activation for point cloud classification. Brighter colors indicate higher activation values. To highlight the most salient activation, we apply a threshold to filter out points with low activation values. This visualization demonstrates that the model learns to focus on the most discriminative foreground object for classifying the point cloud after adding the frozen LLM transformer.
@@
2D VQA.   We investigate the activation of visual tokens in the 2D VQA task in Fig. B. With the METER (Dou et al., 2022) framework initializing the visual backbone from CLIP (Radford et al., 2021) weights, the quality of feature activation is reasonable and mostly concentrates on the target regions for both the baseline METER and our “METER+LLaMA.” However, we can still witness that feature activation better aligns with the images and questions than the noisy attention heads. Furthermore, the activation of “METER+LLaMA” is also slightly advantageous over METER with less scattering, such as better concentrating on the regions of light and leaves at high thresholds.
@@
Figure B: Visualization of attention scores and feature activation for 2D VQA. We are able to visualize attention scores because METER uses $\mathtt{CLS}$ token. Both low and high thresholds for feature activation are displayed to illustrate the concentration on relevant regions.
@@
3D VQA.   We analyze the effect of a frozen LLM transformer for 3D VQA and further confirm our hypothesis. As in Fig. C, we compare the activation of visual tokens, which are seed points in VoteNet (Qi et al., 2019), before and after incorporating the LLaMA transformer. To provide the context, the scenes in SQA3D (Ma et al., 2023) are projected onto the bird’s-eye-view (BEV). From the visualizations, we clearly observe that the feature activation exhibits sharper concentration to the directions guided by language after adding LLaMA, such as the “table behind me” areas in the left figure and “to my left side” areas in the right figure. Therefore, these indicate that the added LLM transformer filters the informative points and augments them for downstream question answering.
@@ A.3 Quantitative Evidence
Pseudo-mask generation.   To generate pseudo masks from feature activation, we are motivated by Fig. 2(a) and treat the highly-activated regions as the final result. To generate pseudo masks with attention scores, we first sum the scores from all the attention heads and follow a similar procedure of treating highly-scored regions as pseudo masks. Concretely, given a feature activation $z$, generating a pseudo-mask is as straightforward as $M_{p}=(z>t)$, where $t$ is a threshold between 0 and 1. Although the process is natural, we notice that selecting the thresholds for activation or score significantly affects the quality of pseudo masks. Therefore, we always automatically choose the threshold maximizing the mIoU between the pseudo and ground truth masks to avoid threshold tuning and enable a fair comparison. The algorithm of choosing the best threshold for the activation or attention scores for each image is as below,
@@
Full results with both magnitude and frequency activation.   Our comparison of mIoU in Fig. 4 only visualizes the larger mIoU from the activation of magnitude or frequency for clarity. To supplement the comprehension, we display the complete statistics in Fig. D. As illustrated, both ViT-S-LLaMA and ViT-B-LLaMA have better pseudo mask quality from feature activation than attention scores and directly support our hypothesis. With the new statics from both activation types, we additionally notice that the neural network layers have varied preferences over magnitudes or frequencies. However, the ViT-LLaMA features still have better fidelity compared to attention scores and features from ViT. As stated in Sec. 6.2, either magnitude or frequency is an intuitive but lossy way to understand feature representation. Thus, future work is needed to further investigate the advantages and properties of individual layers.
Figure D: Visualization of mIoU between the ground truth masks and pseudo masks generated from attention scores/feature activation. This figure supplements Fig. 4 by providing the mIoU for both magnitude and frequency activation, where Fig. 4 selects the better one for the clarity of illustration.
A.4 Implementation Details in Deriving Hypothesis
Magnitude and frequency activation   Our visualization in Sec. 6 and Fig. 2(a) calculates the feature activation on magnitude or frequency domains to reflect their concentration of target objects. We illustrate the Pytorch-style pseudo-code for our operations in Fig. E. For magnitude, we simply compute the L2 norm of features after centering them. Similarly, the activation of the frequency domain is the norm of angle vectors after the 1D Fourier transformation and centering. Finally, the activation values are normalized to 0 and 1 for visualization and quantitative analysis.
@@ B.1 Ablation Study on Design Choices
This section provides supplemental results for the analysis on design choices in Sec. 5.1.
@@
Table A: Ablation studies on model capacity and finetuning.
@@
Results with 100 epochs of training.   We conduct the ablation studies mostly on 100 epochs to balance the computation and fidelity of conclusions. In addition to the experiments lasting 300 epochs in Sec. 5.1, we supplement it with experiments lasting 100 epochs, summarized in Table A. According to the numbers, adding a frozen LLM transformer as a visual encoder is still effective, improving the accuracy of baseline ViT. In addition, we highlight that finetuning is beneficial under insufficient training (100 epochs) but it hurts the accuracy when trained longer due to overfitting, which is analyzed in the next paragraph. In conclusion, both of the experiments validate our design choices and the effectiveness of using pre-trained LLM transformer blocks as encoder layers.
@@
Loss curves for finetuning.   We analyze the design choice of finetuning the LLM transformer in Sec. 5.1. Fig. F below shows the loss curves for both training and validation sets during the training process. The training loss is much larger than the validation loss because their functions are different: the training loss is label-smoothing cross-entropy while the validation loss is the regular cross-entropy loss. With the trend in Fig. F, we conclude that jointly finetuning the pretrained LLM transformer might not benefit the performance while making the training process more complicated. Therefore, our experiments adopt the simple solution of freezing the LLM transformer.
@@
Figure F: Loss curves on the training and validation set for finetuning the LLaMA transformer (ViT-S-LLaMA-FT) or not (ViT-S-LLaMA). We use solid lines to denote training losses and dashed lines to denote validation losses. Though finetuning shows advantages in the beginning, it finally hurts the performance due to overfitting (a larger loss on the validation set compared to not finetuning).
B.2 Breakdown Metrics for 3D VQA

### v2

@@ 6.3 Information Filtering Hypothesis on Other Tasks
The previous sections mainly discuss our information filtering hypothesis in terms of image classification. Meanwhile, we also discover supportive evidence of our hypothesis on various other tasks. This section investigates action recognition as an example, and Sec. A.2 covers additional tasks including point cloud classification, 2D VQA, and 3D VQA.
@@
We mainly analyze the information filtering hypothesis in action recognition qualitatively, because the ground truth of “relevant regions” is difficult to quantify for this task. In practice, we follow a similar procedure in Sec. 6.1 and visualize the activation of video tokens in Fig. 5, and display the most highly activated tokens according to either low or high thresholds for better clarity. At a low activation threshold, we notice that the video tokens from ViT-S-LLaMA better capture the foreground areas of hands and manipulated objects than ViT-S. With the video tokens in VideoMAE activated both spatially and temporally, we further increase the threshold to demonstrate its ability to select informative frames. As in the “high threshold” row of Fig. 5, ViT-S-LLaMA more accurately focuses on the middle frames with actual human-object interaction. Therefore, we conclude that the informative video tokens are indeed distinguished and augmented in action recognition, which aligns with the information filtering hypothesis.
7 Conclusion
@@
In this work, we explore the unexpected capability of large language models (LLMs) as encoders for visual tasks, a significant departure from their conventional text-based applications. By seamlessly integrating a frozen transformer block from pre-trained LLMs into visual encoders, we observe consistent performance enhancements across diverse visual challenges, including 2D image and video classification, 3D point cloud classification, motion forecasting, and 2D and 3D vision-language tasks. This phenomenon, underpinned by our proposed information filtering hypothesis, highlights the inherent adaptability and versatility of LLMs for more general representation learning. We hope that our insights will catalyze further exploration into the uncharted fields of LLM applications and foster innovative strategies to harness their potential in novel ways.
Discussion and limitations.   We have validated the capability of pre-trained, frozen language transformers across a wide spectrum of visual tasks. It is important to note that our goal is to methodically explore this under-investigated problem. Therefore, our experiments are designed to maximize the diversity of tasks under fair comparisons with well-established or competitive baselines, rather than striving for state-of-the-art performance for all tasks, which is also constrained by our computational resources. We leave scaling up the experiments to state-of-the-art levels for all tasks as interesting future work. Meanwhile, we also notice that our information filtering hypothesis has not covered several intriguing questions, e.g., how to quantify the functions of different layers and analyze how the training process facilitates the visual token features to cooperate with the language transformer, which are also meaningful directions.
Acknowledgement.   We thank Baifeng Shi and Shoufa Chen for their valuable discussion and suggestions on implementation and model training. This work was supported in part by NSF Grant 2106825, NIFA Award 2020-67021-32799, the Jump ARCHES endowment through the Health Care Engineering Systems Center, and the IBM-Illinois Discovery Accelerator Institute. This work used NVIDIA GPUs at NCSA Delta through allocation CIS220014 and CIS230012 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by NSF Grants #2138259, #2138286, #2138307, #2137603, and #2138296.
A Thorough Discussion on the Information Filtering Hypothesis
@@ A.1 Detailed Derivation of the Hypothesis
which is identical to Eqn. 3 in the main paper, we further separate the visual tokens into the subset of informative tokens $V_{i}$ and uninformative tokens $V_{u}$. Intuitively, the tokens corresponding to the foreground objects are informative, and the background ones are uninformative. This changes Eqn. A into,
@@
Using the same notation, our observation on the feature activation can be stated as: the attention weights for informative tokens $\{w_{i},i\in V_{i}\}$ are still noisy after incorporating the frozen LLM transformer, while the final visual tokens shown as below have emergent concentration on target regions:
$z_{L}^{2}[i]=\mathbf{F}_{L}^{2}\cdot\mathbf{F}_{LM}^{F}\cdot\mathbf{F}_{LM}^{A}(z_{L}^{1}[i]),\ \text{where}\ i\in V_{i}.$ (C)
Combining Eqn. B and Eqn. C inspires us to express our hypothesis in terms of the connection between visual and $\mathtt{CLS}$ tokens with Eqn. D below, where the added modules $\mathbf{F}_{L}^{2}\cdot\mathbf{F}_{LM}^{F}\cdot\mathbf{F}_{LM}^{A}$ augment the informative tokens $V_{i}$ and lead to better prediction:
@@ A.2 Information Filtering Hypothesis on Other Tasks
This section provides the evidence for our information filtering hypothesis in other tasks, supplementing the discussion on image classification (Sec. 6.1) and action recognition (Sec. 6.3). Specifically, we observe that the frozen LLM transformer selects and amplifies information tokens in point cloud classification, 2D visual question answering (VQA), and 3D VQA. The tasks of motion forecasting and image-text retrieval are not illustrated, because it is more abstract to intuitively define their “informative” tokens. For the investigated tasks, we mainly analyze qualitatively because the ground truth for relevant regions is ambiguous on such tasks, unlike the segmentation masks for image classification (Sec. 6.2).
@@
Point cloud classification.   We visualize the activation of the point tokens in point cloud classification before and after adding the frozen LLM transformer in Fig. A. In the examples of chairs and desks, we observe that “PointBERT-LLaMA” concentrates less on the background (chairs, indicated with red arrows) and more on the actual object surfaces (desks, indicated with red arrows). This demonstrates that the frozen LLM transformer learns to focus on the informative points, which is consistent with our hypothesis.
@@
Figure A: Visualization of feature activation for point cloud classification. Brighter colors indicate higher activation values. To highlight the most salient activation, we apply a threshold to filter out points with low activation values. This visualization demonstrates that the model learns to focus on the most discriminative foreground object for classifying the point cloud after adding the frozen LLM transformer. For visual clarity, we use red arrows to indicate the key regions to observe.
@@
2D VQA.   We investigate the activation of visual tokens in the 2D VQA task in Fig. B. With the METER (Dou et al., 2022) framework initializing the visual backbone from CLIP (Radford et al., 2021) weights, the quality of feature activation is reasonable and it mostly concentrates on the target regions for both the baseline METER and our “METER-LLaMA.” However, we can still witness that feature activation better aligns with the images and questions than the noisy attention heads. Furthermore, the activation of “METER-LLaMA” is also slightly advantageous over METER with less scattering, such as better concentrating on the regions of light and leaves at high thresholds.
@@
Figure B: Visualization of attention scores and feature activation for 2D VQA. We are able to visualize attention scores, because METER uses the $\mathtt{CLS}$ token. Both low and high thresholds for feature activation are displayed to illustrate the concentration on relevant regions.
@@
3D VQA.   We analyze the effect of a frozen LLM transformer for 3D VQA and further confirm our hypothesis. As in Fig. C, we compare the activation of visual tokens, which are seed points in VoteNet (Qi et al., 2019), before and after incorporating the LLaMA transformer. To provide the context, the scenes in SQA3D (Ma et al., 2023) are projected onto the bird’s-eye-view (BEV). From the visualizations, we clearly observe that the feature activation exhibits sharper concentration on the directions guided by language after adding LLaMA, such as the “table behind me” areas in the left figure and “to my left side” areas in the right figure. Therefore, these indicate that the added LLM transformer selects the informative points and augments them for downstream question answering.
@@ A.3 Quantitative Evidence
Pseudo-mask generation.   To generate pseudo-masks from feature activation, we are motivated by Fig. 3(a) and treat the highly-activated regions as the final result. To generate pseudo-masks with attention scores, we first sum the scores from all the attention heads and follow a similar procedure of treating highly-scored regions as pseudo-masks. Concretely, given a feature activation $z$, generating a pseudo-mask is as straightforward as $M_{p}=(z>t)$, where $t$ is a threshold between 0 and 1. Although the process is natural, we notice that selecting the thresholds for activation or score significantly affects the quality of pseudo-masks. Therefore, we always automatically choose the threshold maximizing the mIoU between the pseudo-masks and ground truth masks to avoid threshold tuning and enable a fair comparison. The algorithm of choosing the best threshold for the activation or attention scores for each image is as below,
@@
Full results with both magnitude and frequency activation.   Our comparison of mIoU in Fig. 4 only visualizes the larger mIoU from the activation of magnitude or frequency for clarity. To supplement the comprehension, we display the complete statistics in Fig. D. As illustrated, both ViT-S-LLaMA and ViT-B-LLaMA have better pseudo-mask quality from feature activation than attention scores, which directly supports our hypothesis. With the new statistics from both activation types, we additionally notice that the neural network layers have varied preferences over magnitudes or frequencies. However, the ViT-LLaMA features still have better fidelity compared with attention scores and features from ViT. As stated in Sec. 6.2, either magnitude or frequency is an intuitive but lossy way to understand feature representation. Thus, future work is needed to further investigate the advantages and properties of individual layers.
Figure D: Visualization of mIoU between the ground truth masks and pseudo-masks generated from attention scores/feature activation. This figure supplements Fig. 4 by providing the mIoU for both magnitude and frequency activation, where Fig. 4 selects the better one for the clarity of illustration.
A.4 Discussion and Limitations of the Hypothesis
Perspective of usable information.   We supplement the opinions from Xu et al. (2020) that greatly inspire our investigation of the information filtering hypothesis. Xu et al. (2020) propose that a well-trained neural network layer can be considered as a decipher adding usable information into the features and enabling subsequent modules to better infer the latent information. In our information filtering hypothesis, the incorporated modules are indeed acting as deciphers that enlarge the contribution of usable tokens and benefit downstream predictions.
A.5 Implementation Details in Deriving the Hypothesis
Magnitude and frequency activation   Our visualization in Sec. 6 and Fig. 3(a) calculates the feature activation on magnitude or frequency domains to reflect their concentration on target objects. We illustrate the Pytorch-style pseudo-code for our operations in Fig. E. For magnitude, we simply compute the L2 norm of features after centering them. Similarly, the activation in the frequency domain is the norm of the difference between the angle of a token feature vector and the average angle vector of all the tokens within the image after Fourier transformation. Finally, the activation values are normalized to 0 and 1 for visualization and quantitative analysis.
@@ B.1 Ablation Study on Design Choices
This section provides supplementary results for the analysis on design choices in Sec. 5.1.
@@
Table A: Ablation studies on model capacity and fine-tuning.
@@
Results with 100 epochs of training.   We conduct the ablation studies mostly with 100 epochs to balance the computation and fidelity of conclusions. In addition to the experiments lasting 300 epochs in Sec. 5.1, we supplement them with experiments lasting 100 epochs, summarized in Table A. According to the numbers, adding a frozen LLM transformer as a visual encoder layer is still effective, improving the accuracy of the baseline ViT. In addition, we highlight that fine-tuning is beneficial under insufficient training (100 epochs), but it hurts the accuracy when trained longer due to overfitting, which is analyzed in the next paragraph. In conclusion, both experiments validate our design choices and the effectiveness of using pre-trained LLM transformer blocks as encoder layers.
@@
Loss curves for fine-tuning.   We analyze the design choice of fine-tuning the LLM transformer in Sec. 5.1. Fig. F shows the loss curves for both training and validation sets during the training process. The training loss is much larger than the validation loss, because their loss functions are different: the training loss is the label-smoothing cross-entropy, while the validation loss is the regular cross-entropy loss. With the trend in Fig. F, we conclude that jointly fine-tuning the pre-trained LLM transformer might not benefit the performance, yet making the training process more complicated. Therefore, our experiments adopt the simple solution of freezing the LLM transformer.
@@
Figure F: Loss curves on the training and validation sets for fine-tuning the LLaMA transformer (ViT-S-LLaMA-FT) or not (ViT-S-LLaMA). We use solid lines to denote training losses and dashed lines to denote validation losses. Though fine-tuning shows advantages in the beginning, it finally hurts the performance due to overfitting (a larger loss on the validation set compared with not fine-tuning).
B.2 Ablation Study on Network Capacity Across Diverse Tasks
This section analyzes whether the improvement of our approach is the consequence of the increased capacity and supplements Sec. 5.1. Specifically, we experiment with two sets of additional baselines: (1) using additional MLPs, aligning the number of trainable parameters; (2) using randomly initialized LLM transformer blocks, aligning the total number of parameters, to compare with our approach. Please note that the randomly initialized LLM transformer blocks are trained end-to-end with the visual encoders.
We conduct experiments across all the tasks covered in Sec. 4, as shown in Table B. The results show several important findings validating that our performance gains stem from our method rather than the increased network capacity. More importantly, simply adding MLPs is not a consistently beneficial strategy for all the visual tasks and can be detrimental, resulting in inferior performance on some tasks even compared with the plain baselines. This is because naively adding large MLPs or transformer blocks may lead to challenges in optimization and suffers from a small training dataset compared with LLM pre-training.
(a) Image Classification (ImageNet)
Methods Acc
ViT-S 80.1
+ LLaMA (Ours) 80.7
+ MLP 80.4
+ Random LLM 76.9
(b) Point Cloud Recognition (ScanObjectNN)
Methods BG OBJ T50
PointBert 87.4 88.1 83.1
+ LLaMA (Ours) 88.0 88.5 83.8
+ MLP 86.5 87.3 83.4
+ Random LLM 87.2 88.0 82.6
(c) Action Recognition (SSv2)
Methods Acc1 Acc5
ViT-S 64.7 89.2
+ LLaMA (Ours) 65.9 89.9
+ MLP 63.8 88.9
+ Random LLM 64.1 88.8
(d) Motion Forecasting (Argoverse)
[…]
mmTransformer 0.72 1.10 10.7
+ LLaMA (Ours) 0.71 1.08 10.5
+ MLP 0.74 1.16 11.8
+ Random LLM 0.74 1.15 11.5
(e) 2D Retrieval (Flickr30k)
Methods EM1 EM5 EM10
METER 49.66 80.86 89.48
+ LLaMA (Ours) 50.22 82.26 90.08
+ MLP 49.48 81.12 89.58
+ Random LLM 49.80 81.62 89.72
(f) 3D VQA (SQA3D)
Methods EM1 EM10
SQA3D 47.20 86.82
+ LLaMA (Ours) 48.09 89.03
+ MLP 47.14 88.08
+ Random LLM 47.26 88.46
Table B: Addition comparisons for model capacity. We compare our approach of adding the frozen LLM transformer with adding a randomly initialized MLP (“+MLP”) or LLM blocks (“+Random LLM”) and training end-to-end. The results on diverse tasks uniformly support that our improvement is not merely the result of a larger model capacity. Details are in Sec. B.2.
B.3 Varying LLM Transformer Scales
Model Acc
ViT-S 75.25
+ OPT-125M 71.63
+ OPT-350M 71.56
+ OPT-1.3B 75.62
+ OPT-2.7B 75.74
+ OPT-6.7B 76.29
Table C: Accuracy improves with a larger transformer.
[…]
As indicated by the results in Table C, the benefits of frozen language transformer grow with increasing capacity of OPT transformers. The added transformers enhance the performance only with sufficient sizes (1.3B, 2.7B, 6.7B) and hurt the accuracy when the sizes are small (125M, 350M). Therefore, the phenomenon of LLM transformers enhancing visual tasks only “emerges” at sufficient scales.
B.4 Layers to Insert the Frozen LLM Block(s)
Table D presents an ablation study on different architecture variants, including different places to insert the LLM blocks and the number of LLM blocks. Without loss of generality, we leverage the image classification with ViT-S for our ablation study. From the experiments, we mainly have the following discoveries:
LLM block location.   Inserting the frozen LLaMA block at the beginning or the middle of the visual encoder performs worse than our strategy of inserting LLaMA at the end of the encoder. This supports our design choice and verifies the intuition that LLM blocks are more suitable for high-level semantics instead of low-level visual patterns.
LLM block number.   We also experiment with inserting the last 2 blocks from LLaMA, and find it to be better than our default strategy of using a single LLaMA block. Due to computation constraints, we are unable to extend this to diverse computer vision tasks as in Sec. 4, but this avenue presents an interesting direction and broadening our understanding of LLMs.
Model Configuration Acc Top 1
ViT-S Baseline ViT-S 75.32
+ LLaMA at Tail (Ours) Single LLaMA block at the end of ViT 75.84
+ LLaMA at Middle Single LLaMA block at the middle of ViT 75.55
+ LLaMA at Head Single LLaMA block at the head of ViT 72.66
+ 2 LLaMA Blocks 2 LLaMA blocks at the end of ViT 77.10
Table D: Varying the position and number of LLaMA blocks.
B.5 Breakdown Metrics for 3D VQA